OpenStudy (712noni):

24. Solve ln 4+ ln(3x)=2 A) 1.13 B).18 C)1.41 D).62

OpenStudy (sweetburger):

so you can combine logs \[\ln (a)+\ln (b) = \ln(ab)\]

OpenStudy (712noni):

So then it would look like ln(12x)=2?

OpenStudy (sweetburger):

yes correct

OpenStudy (sweetburger):

now rewrite in in exponential form \[\log_{a} x = y\] \[a^y = x\]

OpenStudy (712noni):

10^2=12?

OpenStudy (712noni):

That's most likely wrong sorry

OpenStudy (mykal):

try raising both sides by e to remove the ln

OpenStudy (712noni):

I don't know how

OpenStudy (mykal):

ok ill show you

OpenStudy (mykal):